# Client-Aided Secure Two-Party Computation of Dynamic Controllers

### v1

@@ 1 Introduction
While these methods have advanced encrypted control in a single-server setting, exploring encrypted control in a multi-party computation setting presents potential advantages in security and performance. A benefit of considering a multi-party setting is that it enables the use of other cryptographic primitives, such as secret sharing [17]. A secret sharing scheme enables the distribution of a secret among a group of parties by splitting the secret into multiple shares. Each party stores a share of the secret, which can be recovered by collecting a certain number of shares. Multi-party computation using secret sharing is generally more efficient and secure than homomorphic encryption.
@@
In this paper, we follow the multi-party computation approach for encrypted control and aim to design a protocol for securely implementing dynamic controllers on two servers. For this purpose, we employ a secret sharing scheme to conceal the controller parameters, states, inputs, and outputs from the servers. We then apply a multiplication protocol based on Beaver triples and a bit truncation protocol for controller computation. The former enables efficient multiplication over secret sharing, and the latter prevents controller state overflow due to recursive multiplication. Using these protocols, we propose an efficient construction of a two-party protocol for computing dynamic controllers.
The main contribution of this study is demonstrating the effectiveness of multi-party computation for encrypted dynamic control. Specifically, our contributions are summarized as follows:
  • We realize secure two-party computation of any linear time-invariant controllers without requiring state decryption or input re-encryption. Compared to previous encrypted dynamic controls in a single-server setting [14, 15, 16], the proposed protocol is simpler and significantly reduces computational complexity.
@@ 2 Problem Setting
We employ a client-aided computation model [23] as a computation model for two-party computation. In this setting, although a client who queries two parties does not interact with them during computation, it provides not only protocol inputs but also auxiliary inputs. The parties use the auxiliary inputs to perform a protocol efficiently and return protocol outputs to the client. Furthermore, throughout this paper, we assume the following standard assumptions for two-party computation.
@@
This assumption implies that the parties follow our intended protocol but independently attempt to learn private information of the client from their received messages.
@@ 4 Secret Sharing and Beaver Triple
This section provides a brief review of secret sharing and its arithmetic, as preliminary information for this study.
@@ 4.1 Secret Sharing
A $t$-out-of-$n$ secret sharing scheme splits a secret into $n$ shares, and the share can be recovered by collecting $t$ shares. In this study, we use a standard $2$-out-of-$2$ secret sharing scheme over $\mathbb{Z}_{q}$ with a prime $q$ defined as follows.
Definition 1 ($2$-out-of-$2$ secret sharing)
A $2$-out-of-$2$ secret sharing scheme over $\mathbb{Z}_{q}$ is a tuple of polynomial-time algorithms $\mathsf{Share}$ and $\mathsf{Reconst}$ such that:
@@ 4.2 Beaver Triple
Under Assumption 3, the $2$-out-of-$2$ secret sharing scheme satisfies perfect security [26]. This security means that parties can learn nothing from their own shares, even if their computational capabilities are unlimited.
@@ 5 Two-Party Computation of Dynamic Controllers
In this section, we propose an implementation of the encoded controller (6) over the two-party configuration shown in Fig. 11(b) using the $2$-out-of-$2$ secret sharing scheme. The computation of (6) involves addition, multiplication, and truncation. The addition and multiplication of shares are supported by the arithmetic of the secret sharing scheme, (P4) and (P5), introduced in the previous section. However, the truncation of shares cannot be computed directly from (P2) to (P5).
@@ 5.1 Truncation Protocol
$\left\lceil\frac{m}{2^{\ell}}\right\rfloor=\frac{m-(m\bmod 2^{\ell})}{2^{\ell}}.$
Moreover, if $q$ is a prime such that $q/2>\absolutevalue*{\lceil m/2^{\ell}\rfloor}$ | and $q/2>2^{\ell}$, it holds that
$\left\lceil\frac{m}{2^{\ell}}\right\rfloor$ $=\quantity(\operatorname{inv}(2^{\ell},q)\cdot 2^{\ell})\left\lceil\frac{m}{2^{\ell}}\right\rfloor\bmod q,$
@@
where $\operatorname{inv}(2^{\ell},q)$ is the modular inverse of $2^{\ell}$ modulo $q$, which can be computed using the extended Euclidean algorithm. Here, let $q$ be the prime modulus of the secret sharing scheme in the previous section, and $\lambda\in\mathbb{N}$ be an appropriately chosen security parameter. Suppose that the bit length of a message $m$ is at most $\kappa=\lfloor\log_{2}q\rfloor-\lambda-1$ and $\ell$ is less than $\kappa$. To simplify our protocol, we assume that $\ell$ is public, which is acceptable because it specifies only the bit length of fractional part. Based on the above observation, the $\mathsf{Trunc}$ protocol shown in Protocol 2 computes the truncation of $\ell$ bits of a message $m\in\mathbb{Z}_{\langle\kappa\rangle}$ over shares.
@@
The security of $\mathsf{Trunc}$ is identical to that of the truncation protocol in [27]. Through protocol execution, $P_{2}$ does not receive any messages from $P_{1}$ and thus learns no information. The only information leaked into $P_{1}$ is $m_{r}=m+2^{\ell}r+r^{\prime}+2^{\ell-1}\bmod q$. Under the conditions of bit lengths of $m$, $r$, and $r^{\prime}$, it holds that $m_{r}=m+r^{\prime\prime}\in\mathbb{Z}_{q}$, where $r^{\prime\prime}=2^{\ell}r+r^{\prime}+2^{\ell-1}$ is uniformly distributed over $\mathbb{Z}_{\langle\kappa+\lambda\rangle}$. Intuitively, $m_{r}$ is a masked message of $m$ by $r^{\prime\prime}$, and thus $P_{1}$ cannot learn any information from $m_{r}$. Indeed, the statistical distance between $m+r^{\prime\prime}$ and $r^{\prime\prime}$, $\frac{1}{2}\sum_{x\in\mathbb{Z}_{q}}\absolutevalue*{\Pr[m+r^{\prime\prime}=x]-\Pr[r^{\prime\prime}=x]}$ |, is less than $2^{-\lambda}$ and so is negligible in $\lambda$. Therefore, the $\mathsf{Trunc}$ protocol achieves statistical security, which is a relaxation of the perfect security. However, statistical security is still valid for unconditional adversaries. This implies that it is stronger than computational security achieved by homomorphic encryption [5, 6, 7, 8, 9]. Note that computational security covers only conditional adversaries whose computational capability is limited to polynomial time [28].
@@ 5.2 Controller Computation Protocol
11:   $P_{i}$: Compute $\llbracket\tilde{x}(t+1)\rrbracket$ and $\llbracket\tilde{u}(t)\rrbracket$ by (12)
12:   $P_{i}$: Send $\llbracket\tilde{u}(t)\rrbracket_{i}$ to the client
@@
$\big{\llbracket}\tilde{A}\big{\rrbracket}$ $\leftarrow\mathsf{Share}\quantity\big(\tilde{A}),$ $\ \big{\llbracket}\tilde{B}\big{\rrbracket}$ $\leftarrow\mathsf{Share}\quantity\big(\tilde{B}),$ $\ \big{\llbracket}\tilde{C}\big{\rrbracket}$ $\leftarrow\mathsf{Share}\quantity\big(\tilde{C}),$ (9)
$\big{\llbracket}\tilde{D}\big{\rrbracket}$ $\leftarrow\mathsf{Share}\quantity\big(\tilde{D}),$ $\ \big{\llbracket}\tilde{x}(0)\big{\rrbracket}$ $\leftarrow\mathsf{Share}\quantity\big(\tilde{x}_{0}),$
@@
$\hat{u}(t)=2^{-2\ell}\mathsf{Reconst}\quantity(\llbracket\tilde{u}(t)\rrbracket).$ (13)
@@
$\tilde{x}(t+1)$ $=A\tilde{x}(t)+B\bar{y}(t)+\delta(t),\quad\tilde{x}(0)=\bar{x}_{0},$
@@
for all $t\in\mathbb{N}_{0}$, where $\tilde{x}(t)=\mathsf{Reconst}\quantity(\llbracket\tilde{x}(t)\rrbracket)$, $\tilde{u}(t)=\mathsf{Reconst}\quantity(\llbracket\tilde{u}(t)\rrbracket)$, $\delta(t)\in\mathbb{R}^{n}$ is an error such that $\norm{\delta(t)}_{\infty}\leq 3/2$, and $c\geq 1$ and $\gamma\in(0,1)$ are constants such that $\norm*{\Phi^{t}}\leq c\gamma^{t}$ for all $t\in\mathbb{N}_{0}$.
@@
Lemma 5.3 implies that the state and input computed by the proposed controller (12) are approximations of those of the encoded controller (6) if the bit length of modulus $q$ is larger than $k+\lambda+2+\lfloor\log_{2}\max\{n,p\}\alpha\beta\frac{c}{1-\gamma}\rfloor$ bits. Based on this result, the following theorem states that the control performance of the proposed controller can be arbitrarily close to that of the original controller (2) by choosing the bit length of the fractional part $\ell$ to be sufficiently large.
@@
Protocol 3 sequentially invokes the $\mathsf{Mult}$ and $\mathsf{Trunc}$ protocols as subprotocols in the online phase. In this case, the modular composition theorem in [29] guarantees that the security of the entire protocol is equivalent to the weakest security of the subprotocols. Therefore, the proposed protocol achieves statistical security with respect to $\lambda$.
@@ 5.3.1 LWE-based encrypted controller in [14]
On the server, the encrypted controller requires $n(n+2m+p)+mp$ multiplications between the GSW and LWE ciphertexts and $n(n+2m+p-4)+m(p-1)$ additions of the LWE ciphertexts. According to [16], each GSW-LWE multiplication comprises $\zeta(N+1)^{2}$ scalar multiplications. In addition, each LWE addition is an $(N+1)$-dimensional vector addition. Consequently, the computational complexity of the server is $O((n+m)(n+p)\zeta N^{2})$.
@@ 5.3.2 RLWE-based encrypted controller in [15]
The encrypted controller requires $n+1$ multiplications with linearization, $n+m+p+2$ additions, and $m+p-1$ rotations of the RLWE ciphertexts. Note that the multiplication in the RLWE encryption increases the ciphertext size from $R_{q}^{2}$ to $R_{q}^{3}$. The linearization reduces it to $R_{q}^{2}$, and the rotation replaces the position of vector elements packed in a ciphertext. Each multiplication and addition process includes four multiplications and two additions in $N$-dimensional polynomials, respectively. The linearization and rotation carry out $2\zeta$ $N$-dimensional polynomial multiplications [16], where $\zeta$ is as in the LWE-based scheme. Therefore, the computational complexity of the server is $O((n+m+p)\zeta N\log N)$.
@@ 5.3.4 Proposed protocol (Protocol 3)
One might notice that the computational complexity of the client in Protocol 3 is the same as that of the unencrypted case. This is because the client needs to generate the auxiliary inputs $\mathsf{aux}_{i}$, of which size is $O((n+m)(n+p))$. However, the auxiliary inputs are independent of the protocol inputs (e.g., the plant output $\hat{y}(t)$). This implies that their generation can be addressed in the offline phase if the control system terminates in a finite time. Even if this is not the case, it can be outsourced to a third party called a trusted dealer, who is non-colluding with the parties. Then, the computational complexity can be reduced to $O(m+p)$.
@@
Another direction for reducing the computational complexity is to design a protocol in which the parties emulate the trusted dealer. This is a common problem in multi-party computation. In the next section, we will introduce a pseudorandom correlation function, which is one of the most recent approaches for solving this problem.
@@ 6.1 Brunovsky normal form
$\big{\llbracket}\tilde{x}_{n_{i}}(t+1)\big{\rrbracket}$ $\leftarrow\mathsf{Trunc}\quantity(\big{\llbracket}\tilde{A}_{n_{i}}\big{\rrbracket}\big{\llbracket}\tilde{x}_{n_{i}}(t)\big{\rrbracket}+\big{\llbracket}\tilde{B}_{n_{i}}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},\ell),$ (17)
@@ 7 Numerical Examples
[…]
@@
[…]
@@
[…]

### v2

@@ 1 Introduction
While these methods have advanced encrypted control in a single-server setting, exploring encrypted control in a multi-party computation setting presents potential advantages. Multi-party computation using secret sharing [17] is generally more computationally efficient and secure than homomorphic encryption in exchange for an increase in communication costs. Furthermore, multi-party computation offers the ability to evaluate various functionalities. This property motivates the implementation of encrypted control in a multi-party setting because homomorphic encryption typically requires substantial computational resources for computing mathematical operations beyond addition and multiplication.
@@
In this paper, we aim to design an efficient protocol for securely implementing dynamic controllers. For this purpose, we follow the multi-party computation approach and perform bit truncation over secret sharing to prevent controller state overflow due to recursive multiplication. The main contribution of this study is to demonstrate the effectiveness of multi-party computation for encrypted dynamic control. Our contributions are summarized as follows:
  • We realize secure two-party computation of any linear time-invariant controllers without requiring state decryption or input re-encryption. This requirement of previous encrypted dynamic controls in a single-server setting [14, 15, 16] is undesirable because of increasing computation and communication efforts on a client. Compared to the previous schemes, the proposed protocol is simpler and significantly reduces the computation costs of a client.
@@ 2 Problem Setting
We adopt a client-aided computation model [23] for two-party computation. In this setting, the client who initiates queries to two parties provides not only the protocol inputs but also auxiliary inputs, although it does not interact with them during the computation. The parties use these auxiliary inputs to execute the protocol efficiently and return the outputs to the client. We also assume peer-to-peer secure channels as a network model, which can be realized by standard symmetric-key encryption. Furthermore, throughout this paper, we assume the following standard assumptions for two-party computation.
@@
This assumption implies that the parties follow our intended protocol but independently attempt to learn private information of the client from their received messages. It also prohibits external adversaries from corrupting or accessing both parties.
@@ 4 Secret Sharing and Beaver Triple
This section provides a brief review of secret sharing and its arithmetic [17, 25].
@@ 4.1 Secret Sharing
A $t$-out-of-$n$ secret sharing scheme splits a secret into $n$ shares, and the share can be recovered by collecting $t$ shares. In this study, we use a standard 22-out-of-22 secret sharing scheme over $\mathbb{Z}_{q}$ with a prime $q$ defined as follows.
Definition 1 (22-out-of-22 secret sharing)
A 22-out-of-22 secret sharing scheme over $\mathbb{Z}_{q}$ is a tuple of polynomial-time algorithms $\mathsf{Share}$ and $\mathsf{Reconst}$ such that:
@@ 4.2 Beaver Triple
Under Assumption 3, the 22-out-of-22 secret sharing scheme satisfies perfect security [26]. This security means that parties can learn nothing from their own shares, even if their computational capabilities are unlimited.
@@ 5 Two-Party Computation of Dynamic Controllers
In this section, we propose an implementation of the encoded controller (6) over the two-party configuration shown in Fig. 11(b) using the 22-out-of-22 secret sharing scheme. The computation of (6) involves addition, multiplication, and truncation. The addition and multiplication of shares are supported by the arithmetic of the secret sharing scheme, (P4) and (P5), introduced in the previous section. However, the truncation of shares cannot be computed directly from (P2) to (P5).
@@ 5.1 Truncation Protocol
$\left\lceil\frac{m}{2^{\ell}}\right\rfloor$ $=\frac{m-(m\bmod 2^{\ell})}{2^{\ell}},$
@@
if $q$ is a prime such that $|q/2>\absolutevalue*{\lceil m/2^{\ell}\rfloor}$ and $q/2>2^{\ell}$. Notice that $\operatorname{inv}(2^{\ell},q)$ is the modular inverse of $2^{\ell}$ modulo $q$ and can be computed using the extended Euclidean algorithm. Here, let $q$ be the prime modulus of the secret sharing scheme in the previous section, and $\lambda\in\mathbb{N}$ be an appropriately chosen security parameter. Suppose that the bit length of a message $m$ is at most $\kappa=\lfloor\log_{2}q\rfloor-\lambda-1$ and $\ell$ is less than $\kappa$. To simplify our protocol, we assume that $\ell$ is public, which is acceptable because it specifies only the bit length of fractional part. Based on the above observation, the $\mathsf{Trunc}$ protocol shown in Protocol 2 computes the truncation of $\ell$ bits of a message $m\in\mathbb{Z}_{\langle\kappa\rangle}$ over shares.
@@
The security of $\mathsf{Trunc}$ is identical to that of the truncation protocol in [27]. Through protocol execution, $P_{2}$ does not receive any messages from $P_{1}$ and thus learns no information. The only information leaked into $P_{1}$ is $m_{r}=m+2^{\ell}r+r^{\prime}+2^{\ell-1}\bmod q$. Under the conditions of bit lengths of $m$, $r$, and $r^{\prime}$, it holds that $m_{r}=m+r^{\prime\prime}\in\mathbb{Z}_{q}$, where $r^{\prime\prime}=2^{\ell}r+r^{\prime}+2^{\ell-1}$ is uniformly distributed over $\mathbb{Z}_{\langle\kappa+\lambda\rangle}$. Intuitively, $m_{r}$ is a masked message of $m$ by $r^{\prime\prime}$, and thus $P_{1}$ cannot learn any information from $m_{r}$. Indeed, the statistical distance between $m+r^{\prime\prime}$ and $r^{\prime\prime}$, $|\frac{1}{2}\sum_{x\in\mathbb{Z}_{q}}\absolutevalue*{\Pr[m+r^{\prime\prime}=x]-\Pr[r^{\prime\prime}=x]}$, is less than $2^{-\lambda}$ and so is negligible in $\lambda$. Therefore, the $\mathsf{Trunc}$ protocol achieves statistical security, which is a relaxation of the perfect security. However, statistical security is still valid for unconditional adversaries. This implies that it is stronger than computational security achieved by homomorphic encryption [5, 6, 7, 8, 9]. Note that computational security covers only conditional adversaries whose computational capability is limited to polynomial time [28].
@@ 5.2 Controller Computation Protocol
[…]
@@
[…]
@@
$\hat{u}(t)=2^{-2\ell}\tilde{u}(t),\quad\tilde{u}(t)=\mathsf{Reconst}\quantity(\llbracket\tilde{u}(t)\rrbracket).$ (13)
@@
$\tilde{x}(t+1)$ $=\left\lceil\quantity(\bar{A}\tilde{x}(t)+\bar{B}\bar{y}(t))/2^{\ell}\right\rfloor+w(t),\quad\tilde{x}(0)=\bar{x}_{0},$
@@
for all $t\in\mathbb{N}_{0}$, where $\tilde{x}(t)=\mathsf{Reconst}\quantity(\llbracket\tilde{x}(t)\rrbracket)$, $\tilde{u}(t)=\mathsf{Reconst}\quantity(\llbracket\tilde{u}(t)\rrbracket)$, $w(t)\in\{-1,0,1\}^{n}$, and $c\geq 1$ and $\gamma\in(0,1)$ are constants such that $\norm*{\Phi^{t}}\leq c\gamma^{t}$ for all $t\in\mathbb{N}_{0}$.
@@
Lemma 5.3 implies that the reconstruction of the proposed controller (12) is equivalent to the encoded controller (6) with the error $w(t)$ induced by the $\mathsf{Trunc}$ protocol. Based on this result, the following theorem states that the control performance of the proposed controller can be arbitrarily close to that of the original controller (2) by choosing the bit length of the fractional part $\ell$ to be sufficiently large.
@@
Note that the term $k+\lambda+2$ in (14) is essential to guarantee the correctness and security of the $\mathsf{Trunc}$ protocol. The last term of (14) is induced by encoding and mainly depends on $2^{\ell}$ in $\beta$. This implies a trade-off between communication costs and precision. This trade-off can be improved by optimizing the lower bound of $\ell$ in (15).
@@
During the execution of Protocol 3, each party $P_{i}$ receives the $i$th shares in (9), (10), and (11) from the client and can access public parameters $k$, $\ell$, $q$, and $\lambda$. It also obtains the masked messages $(d,e)$ and $m_{r}$ (if $i=1$) for each invocation of $\mathsf{Mult}$ and $\mathsf{Trunc}$, respectively. Here, the subprotocols are sequentially invoked in the online phase. In this case, the modular composition theorem in [29] guarantees that the security of the entire protocol is equivalent to the weakest security of the subprotocols. Therefore, the proposed protocol achieves statistical security with respect to $\lambda$.
@@ 5.3.1 LWE-based encrypted controller in [14]
On the server, the encrypted controller requires $n(n+2m+p)+mp$ multiplications between the GSW and LWE ciphertexts and n(n+2m+p4)+m(p1)n(n+2m+p-4)+m(p-1) additions of the LWE ciphertexts. According to [16], each GSW-LWE multiplication comprises $\zeta(N+1)^{2}$ scalar multiplications. In addition, each LWE addition is an $(N+1)$-dimensional vector addition. Consequently, the computational complexity of the server is $O((n+m)(n+p)\zeta N^{2})$.
@@ 5.3.2 RLWE-based encrypted controller in [15]
The encrypted controller requires $n+1$ multiplications with linearization, $n+m+p+2$ additions, and m+p1m+p-1 rotations of the RLWE ciphertexts. Note that the multiplication in the RLWE encryption increases the ciphertext size from $R_{q}^{2}$ to $R_{q}^{3}$. The linearization reduces it to $R_{q}^{2}$, and the rotation replaces the position of vector elements packed in a ciphertext. Each multiplication and addition process includes four multiplications and two additions in $N$-dimensional polynomials, respectively. The linearization and rotation carry out $2\zeta$ NN-dimensional polynomial multiplications [16], where $\zeta$ is as in the LWE-based scheme. Therefore, the computational complexity of the server is $O((n+m+p)\zeta N\log N)$.
@@ 5.3.4 Proposed protocol (Protocol 3)
One might notice that the computational complexity of the client in Protocol 3 is the same as that of the unencrypted case. This is because the client needs to generate the auxiliary inputs $\mathsf{aux}_{i}$, of which size is $O((n+m)(n+p))$. However, the auxiliary inputs are independent of the protocol inputs (e.g., the plant output $\hat{y}(t)$). This implies that their generation can be addressed in the offline phase if the control system terminates in a finite time. Even if this is not the case, it can be outsourced to a third party called a trusted dealer, who is non-colluding with the parties. Then, the computational complexity of the client can be reduced to $O(m+p)$.
@@
Protocol 3 can be extended to $M$ parties using an $M$-out-of-$M$ secret sharing scheme. Then, the complexities in Table 1 increase linearly with $M$. Meanwhile, Assumption 3 can be relaxed to allow the collusion of any M1M-1 parties, which mitigates security risks. Recall that the proposed scheme is compromised when all parties are corrupted, unlike the single-server setting using homomorphic encryption.
@@ 6.1 Brunovsky normal form
$\big{\llbracket}\tilde{x}_{n_{i}}(t+1)\big{\rrbracket}$ $\leftarrow\mathsf{Trunc}\quantity(\big{\llbracket}\tilde{A}_{n_{i}}\big{\rrbracket}\big{\llbracket}\tilde{x}_{n_{i}}(t)\big{\rrbracket}+\big{\llbracket}\tilde{B}_{n_{i}}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},\ell),$
@@ 7 Numerical Examples
Consider a PID controller of parallel form in [14]. The controller parameters of (2) are $A=\quantity[\begin{smallmatrix}2-N_{d}&N_{d}-1\\ 1&0\end{smallmatrix}]$, $B=[1\ 0]^{\top}$, $C=[c_{1}\ c_{2}]$, and $D=d$, where $c_{1}=K_{i}T_{s}-N_{d}^{2}K_{d}/T_{s}$, $c_{2}=N_{d}K_{i}T_{s}-K_{i}T_{s}+N_{d}^{2}K_{d}/T_{s}$, and $d=K_{p}+N_{d}K_{d}/T_{s}$. Here, $K_{p}$, $K_{i}$, and $K_{d}$ are proportional, integral, and derivative gains, respectively, $N_{d}\in\mathbb{N}$ is the parameter of a derivative filter, and $T_{s}\in\mathbb{R}$ is a sampling time. Note that, if $2-N_{d},N_{d}-1\in\mathbb{Z}_{\langle k-\ell\rangle}$, the parameters $A$ and $B$ are already integers. Then, the proposed controller (12) can be simplified as
@@
Suppose that the plant (1) and sampling time are given by the single-input and single-output system in  [38, Eq. (2), with $\alpha=0.2$] and $T_{s}=0.1$ s, respectively. Let $\epsilon=2^{-10}$, $\lambda=80$ bit, $k-\ell=8$ bit, $x_{p,0}=[100\ 100\ 100\ 100]^{\top}$, and $x_{0}=[0\ 0]^{\top}$. For instance, the PID controller with the parameters $A=\quantity[\begin{smallmatrix}1&0\\ 1&0\end{smallmatrix}]$, $B=[1\ 0]^{\top}$, C=[2.73689272.96540833]C=[2.7368927\ {-2.96540833}], and D=5.01071167D=-5.01071167 satisfies Assumption 1, Assumption 2, and (15) with $\ell=32,40,48,56$ bit. The bit length of modulus $q$ can be chosen as $256$ bit to satisfy (14) for all the choices of $\ell$. Fig. 3 shows the input errors $\norm{u(t)-\hat{u}(t)}$. In this figure, the input errors remain below $\epsilon=2^{-10}$ for all $\ell=32,40,58,64$ and $t=0,\dots,50$, thereby supporting Theorem 5.5. Furthermore, the figure suggests that increasing the bit length $\ell$ can reduce the errors.
@@
Suppose the plant is given as the four-tank process in [39] with $h_{1}^{0}=12.4$ cm, $h_{2}^{0}=12.7$ cm, $h_{3}^{0}=1.8$ cm, $h_{4}^{0}=1.4$ cm, $v_{1}^{0}=3$ V, $v_{2}^{0}=3$ V, $\gamma_{1}=0.7$, $\gamma_{2}=0.6$ and $T_{s}=0.5$ s. Let $\epsilon$, $\lambda$, $k$-$\ell$, and $\ell$ be as in the PID control example. Let $x_{p,0}=[10\ 10\ 10\ 10]^{\top}$ and $x_{0}=[0\ 0\ 0\ 0]^{\top}$. The controller with the parameters,